The sum of two numbers is 148.The larger number is 16 more than twice the smaller number.What are the numbers?
Found 2 solutions by azhou4, stanbon:
Answer by azhou4(12)

In this problem, each sentence is an equation. The sum of two numbers is 148 means you add the two numbers together. Let x and y be the two unknown numbers.
So x+y=148
The larger number is 16 more than twice the smaller number
y=16+2x
SO your two equations are
x+y=148
y=16+2x
You can plug the second equation into the first equation.
x+16+2x=148
3x+16=148
3x=148-16
3x=132
x=132/3
x=44
So now you know x=44. Just plug this into the first or second equation and you have your answer! Hope this helps!

Answer by stanbon(75887)

The sum of two numbers is 148.The larger number is 16 more than twice the smaller number.What are the numbers?
---------------------------
Using one variable:
1st number: x
2nd number 148-x
-----------------------
Equation:
You really don't know which number is the larger so pick the most
convenient in the equation:
148-x = 2x+16
3x = 132
x = 44
---
148 - x = 104
===================
Cheers,
Stan H.